A copper wire of length 1 meter, cross sectional area 7.1 X 10‑6 m2 and Young’s modulus 11 X 1010 N/m2 has a mass load hung on it. If its increase in length is 1 mm, what is the value of the mass? a) 40kg b) 80kg c) 200kg d) 400kg e) 800kg Please help me. Thank you
Answer #1

se the formula for Young's modulus, E

E = (F/Ao)/(dL/Lo) = (F * Lo)/(Ao * dL)

E - young's modulus = 11 * 10^10 N/m2
F - force applied = 200kg * 9.8 m/s2
Lo = original length of object = 2m
Ao = cross-sectional area through which force is applied = 7.1 * 10^-6 m2
dL = amount length of object changes - what u r solving for

rearrange above to solve for dL

dL = (F * Lo)/(Ao * E) =(mg * Lo)/(Ao * E)
dL*(Ao * E)/Lo*g=m

1x10^-3m*(7.1 X 10‑6 m2* 11 X 10^10 N/m2)1*9.8=80Kg

option B